Nevey:

There are 7 tigers at the zoo. Each year, the number of tigers increases by 10%. There are 15 eagles at the zoo. Each year, 2 new eagles join the zoo. Part A: Write functions to represent the number of tigers and eagles at the zoo throughout the years. (4 points) Part B: How many tigers are at the zoo after 10 years? How many eagles are at the zoo after the same number of years? (2 points) Part C: After approximately how many years is the number of tigers and eagles the same? Justify your answer mathematically. (4 points)

dude:

Have you written anything on this before I write anything? xd

Nevey:

no i havent.

dude:

Do you know the general exponential formula?

Nevey:

no

dude:

Since the population of tigers is increased you want to do exponential growth \(y = a(1 + r)^x\)

dude:

Does that seem familiar to you?

Nevey:

kinda but not really

Nevey:

@dude

dude:

a is the original amount and the r is the rate at which it increases Turn 10% into a decimal, do you know how to do that?

Nevey:

its 0.1right?

dude:

Yes

Nevey:

alright

dude:

\(y = a(1 + r)^x\) a = original r = % as decimal Substitute what we have

Nevey:

y=7(1+0.1)^x

Nevey:

thats part of part a right?

dude:

Ya

Nevey:

ok now what?

dude:

To get part B Substitute 10 in for x x = years \(y=7(1+0.1)^x\)

dude:

For part C you want to find the equation for eagles at the zoo This is a linear growth \(y=mx+b\) m is the rate b is the original amount

dude:

`There are 15 eagles at the zoo. Each year, 2 new eagles join the zoo.`

dude:

From there you want to set both equations equal to each other \(y=7(1+0.1)^x\) and (the equation you get from above)

Nevey:

y = 2x + 15 y=7(1+0.1)^10 @dude do i just stop there?

1 attachment
dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude From there you want to set both equations equal to each other \(y=7(1+0.1)^x\) and (the equation you get from above) \(\color{#0cbb34}{\text{End of Quote}}\)

Nevey:

I know, i read that, but is that not what i did :/ im confused. what do i do after putting one equation above the other?

dude:

Set them equal \(7(1+0.1)^x=2x+15\) Then solve for x

Nevey:

well, the first equation is equal to 17. i did that in part B. so i just solve the equation to find how many eagles right?

dude:

No no you’re solving for them individually, you want to get the x on the same side

Nevey:

oh wait lol now i get it. ok sorry :)

Nevey:

@dude this is what i got. is it right?

1 attachment
Nevey:

well i have to turn it in now. thank you so much for your help. i will just hope its all correct. :)